# AmodalGen3D: Generative Amodal 3D Object Reconstruction from Sparse Unposed Views

### v1

@@ B Additional Results
In this section, we present more results on the 3D amodal reconstruction results at single-object level in Sec. B.1 (additional results on the GSO dataset); we also give more in-the-wild and in-the-scene results in Sec. B.2.
For the efficiency evaluation, our AmodalGen3D can typically generate a single 3D object within 15 seconds, which causes a slight computation overhead compared to TRELLIS [xiang2025structured] and Amodal3R [wu2025amodal3r]. However, our method leads to much better reconstruction results, which is worth the increased inference time.
B.1 Additional Single-Object Results
Fig. 11 and Fig. 12 show more examples on the single-object level reconstruction results. We present results with 1, 2, and 4 views, validating the effectiveness of our method. We still compare with TRELLIS [xiang2025structured] and Amodal3R [xu2024amodal] in this part to show that our method can fix inconsistencies and infer high-quality unobserved structures.
B.2 Additional In-the-wild & In-the-scene Results
We give more in-the-wild and in-the-scene examples in Fig. 13 and Fig. 14, showcasing that AmodalGen3D achieves superior fidelity and completeness under various occlusion conditions in real-world scenarios.
Figure 11: More qualitative results with different sparse-view settings (1, 2, and 4 views). We conduct comparisons with TRELLIS and Amodal3R on the Google Scanned Object dataset, demonstrating the effectiveness of our method.
Figure 12: We present more examples generated by AmodalGen3D on the GSO dataset.
Figure 13: Additional in-the-wild amodal 3D object reconstruction results on the MipNeRF 360 [barron2022mipnerf360] dataset.
Figure 14: Additional in-the-scene amodal 3D object reconstruction results on the Hypersim [roberts2021hypersim] dataset.

### v2

@@ B Additional Results
In this section, we present more qualitative results on the 3D amodal generation results at the single-object level in Sec. B.2 and an occlusion degree analysis in Sec. B.1 (additional results on the GSO dataset); we also give more in-the-wild and in-the-scene results in Sec. B.3.
For the efficiency evaluation, our GENA3D can typically generate a single 3D object within 15 seconds, which causes a slight computation overhead compared to TRELLIS [xiang2025structured] and Amodal3R [wu2025amodal3r]. However, our method leads to much better generation results, which is worth the increased inference time.
B.1 Occlusion Degree
We randomly select 500 objects from the GSO dataset and render 3D-consistent masks at occlusion levels of 20%, 50%, and 75%, as suggested by Reviewer ruEG. We evaluate FID, KID, and CLIP-Score using one and two input views (Tab. B.1), demonstrating that our model remains robust under varying degrees of occlusion.
Table B.1: The performance of our GENA3D under different occlusion degrees. We test under the setting of single view and 2 views, with occlusion degrees of 20%, 50%, and 70%.
Occ. Degree View Num. FID $\downarrow$ KID(%) $\downarrow$ CLIP(%) $\uparrow$
20% 1 view 32.28 0.44 82.40
50% 34.16 0.46 82.08
70% 37.22 0.47 81.44
20% 2 views 31.98 0.44 82.54
50% 33.42 0.46 82.32
70% 37.16 0.47 81.66
B.2 Additional Single-Object Results
Fig. B.2 and Fig. B.3 show more examples on the single-object level generation results. We present results with 1, 2, and 4 views, validating the effectiveness of our method. We still compare with TRELLIS [xiang2025structured] and Amodal3R [xu2024amodal] in this part to show that our method can fix inconsistencies and infer high-quality unobserved structures.
B.3 Additional In-the-wild & In-the-scene Results
We give more in-the-wild and in-the-scene examples in Fig. B.4 and Fig. B.5, showcasing that GENA3D achieves superior fidelity and completeness under various occlusion conditions in real-world scenarios.
B.4 Additional comparison with SAM3D
Additionally, we provide more comparisons between our GENA3D and SAM3D in Fig. B.1. As shown in the figure, SAM3D does not generate 3D objects that are fully respective to the input observations. And even with massive training data, the generative ability is still not plausible enough. These comparisons demonstrate the importance of our design of bridging 2D priors and 3D coherence.
Figure B.1: We present more comparisons with SAM3D.
Figure B.2: More qualitative results with different sparse-view settings (1, 2, and 4 views). We conduct comparisons with TRELLIS and Amodal3R on the Google Scanned Object dataset, demonstrating the effectiveness of our method.
Figure B.3: We present more examples generated by GENA3D on the GSO dataset.
Figure B.4: Additional in-the-wild amodal 3D object generation results on the MipNeRF 360 [barron2022mipnerf360] dataset.
Figure B.5: Additional in-the-scene amodal 3D generation results on the Hypersim [roberts2021hypersim] dataset.